# Absolute Calibration of Cluster Mira Variables to Provide a New Anchor for the Hubble Constant Determination

### v1

@@ 4.2 The global slope and zero-points of anchor PLRs
The universality of the slope of Mira PLR in F160W filters is the basic assumption for the distance scale application. C. D. Huang et al. (2018) used two different slopes of mag/dex based on W. Yuan et al. (2017) LMC sample, and mag/dex based on the transformed HST F160W filter. The slope of the globular cluster and LMC PLRs in the /F160W filters are significantly different, presumably due to different period distribution of Miras in these systems. With three independent anchors available, we can determine a global slope of Mira PLRs in calibrated F160W magnitudes. For the LMC, we adopt the precise geometric distance to the LMC (G. Pietrzyński et al., 2019). For the NGC 4258, a geometric maser distance of mag (D. W. Pesce et al., 2020) was adopted that is precise to .
@@
To account for the possible metallicity effects, we adopt a two-parameter solution simultaneously deriving a zero-point and a metallicity coefficient. For the cluster Miras, we derive a mean metallicity of with a scatter of dex after excluding the most metal-poor cluster. For the LMC, a metallicity of dex is taken based on age-metallicity relations for intermediate-age stellar populations in the field and cluster stars (W. Narloch et al., 2023). M. W. Feast (1996) also suggested [Fe/H] dex for Miras in the LMC based on a comparison with cluster variables. Given the lack of metallicity measurements for stellar populations in NGC 4258, we adopt a value of dex based on HII region emission line flux measurements from W. Yuan et al. (2022). This value is consistent with stellar metallicity of blue supergiants (R.-P. Kudritzki et al., 2024) and has been used in the Cepheid-based distance ladder (A. G. Riess et al., 2022a). While the adopted metallicity is appropriate for younger stellar populations, considering no significant metallicity gradient in NGC 4258 (W. Yuan et al., 2022; R.-P. Kudritzki et al., 2024) and the lack of stellar metallicity measurements, this approximation is adopted with a relatively larger uncertainty of 0.15 dex. For the MW field stars, we assume solar metallicity for Miras, but this was not included in the quantitative analysis. Using the aforementioned fixed slope, we estimated the following global solution:
where the mean metallicity of the anchor sample is [Fe/H] dex. We find a small positive metallicity coefficient such that metal-rich Miras are fainter. However, this quantification of metallicity effect is based on our approximation of mean metallicities of Miras in the LMC, in globular clusters and NGC 4258. Fig. 6 displays this linear fit to the zero-points of three anchor galaxies as a function of their adopted metallicities. We find a small variation in the zero-point of mag with or without metallicity term, further suggesting a small effect in the luminosity due to variations in iron abundance.
Furthermore, no metallicity measurements of Miras or intermediate-age populations are available in SNe Ia host galaxies. C. D. Huang et al. (2024) used oxygen-abundance gradient based-on HII region to derive an approximate metallicity range for M101 Miras that falls between solar and LMC values. We assume that the SNe Ia host galaxies have metallicities similar to NGC 4258 with a possible difference of dex, which may increase up to dex with respect to LMC Miras. Given the approximations involved in determining the metallicity coefficient, we do not include this term in the subsequent analysis, but add a systematic uncertainty term accounting for the difference in mean metallicities between anchor and host galaxies.
@@ 4.3 Distances to Type Ia Supernovae host galaxies
The second step of the distance ladder involves using primary calibration of stellar standard candles to determine distances to SNe Ia host galaxies and calibrate peak fiducial luminosity of a SNe Ia. For the Mira distance scale, there are two supernovae host galaxies observed with HST — SN 2005df in NGC 1559 (C. D. Huang et al., 2020) and SN 2011fe in M101 (C. D. Huang et al., 2024). We used a sample of 110 Miras in NGC 1559 and 211 Miras in M101 from aforementioned studies along with their periods and calibrated F160W magnitudes including corrections due to crowding effects. We applied corrections for the Milky Way foreground extinction as described in C. D. Huang et al. (2024).
@@ 4.4 Hubble flow, , and systematics
Table 6 provides a detailed breakdown of the error budget on the . The total uncertainty is dominated by the statistical uncertainty on the fiducial peak luminosity of SNe Ia in two host galaxies. With a global slope that is determined using Miras in three-anchor galaxies, we obtained an improved constrain on the anchor zero-point luminosity of a Mira with d period. The absolute value of the anchor zero-point is similar to C. D. Huang et al. (2024) value, but the uncertainty is reduced significantly considering that the metallicity systematics are also included in the anchor zero-point. The metallicity systematic is propagated through the coefficient of with a plausible difference in the mean metallicity of dex between NGC 4258/Cluster/LMC and SNe Ia host galaxies. We also include a systematic error of anchor distances and also add 0.02 mag for the to F160W transformations for the ground-based LMC and Cluster calibrations. The statistical uncertainties due to SNe Ia host galaxies and the slope of the PLR are similar to those from C. D. Huang et al. (2024). The uncertainty due to interstellar extinction in SNe Ia host galaxies is propagated assuming that the reddening is similar to the LMC, where a difference of 0.02 mag is obtained with and without interstellar reddening correction. The global slope together with improved anchor zero-points provide better constrain on the for individual SNe Ia host galaxies. The difference in of mag between two SNe host galaxies suggests excellent consistency in the fiducial peak luminosity of their SNe Ia. Finally, the inverse-weighted average value of fiducial peak luminosity of SNe Ia is better constrained with an error smaller than the estimated uncertainty in C. D. Huang et al. (2024). These small improvements resulting from additional anchor in the solution provide a significant reduction in the error budget to in the Mira distance ladder.
@@ 5 Discussion
Another approach of quantifying metallicity effects involves comparing the calibrated zero-points of the PLRs in anchor galaxies with different metallicities. Using cluster, LMC, and NGC 4258 Miras, we find a small slope ( mag/dex) of the variation in the zero-points as a function of metallicity. This determination of metallicity effects is also very sensitive to the uncertainties on distances to these systems, as well as the adopted mean-metallicities for their Mira stars. For the distant systems observed with HST, we assume the mean-metallicity based on HII regions that is used in the Cepheid distance ladder (A. G. Riess et al., 2022b). This approximation may not be appropriate for intermediate-age Mira populations, and therefore, we do not include metallicity term in the distance scale analysis. However, these mean metallicity can be used to estimate relative difference in abundances between anchor and host galaxies, assuming that is similar for young and intermediate-age populations. This allowed us to provide a reasonable estimate of systematic uncertainty due to metallicity in the determinations. If we propagate a metallicity term of mag/dex in the Mira distance ladder adopting the metallicity of SNe Ia host galaxies from A. G. Riess et al. (2022a), we find a small difference of units in . Homogeneous spectroscopic metallicity determinations of intermediate-age stars or AGB stars in these galaxies will be useful to precise quantify the metallicity effects based on zero-points of anchor PLRs.
@@
Finally, the Mira based km s-1 Mpc-1 derived in this work provides an independent comparison to Cepheid and TRGB based determinations. As previously mentioned, the uncertainty on the value primarily comes from the peak fiducial luminosity of SNe Ia, and additional SNe Ia hosts with Mira variables are required to minimize these statistical uncertainties. Nevertheless, our accurate determination of is in excellent agreement with the three-anchor baseline solution from Cepheid based local measurements. The value based on intermediate-age Mira stars is also in agreement with population II TRGB based measurements from W. L. Freedman et al. (2025). Higher-precision Mira based values may provide an insight into possible systematics due to stellar standard candles of different age and metallicity populations. Currently, our estimates are in excellent agreement with Cepheid-based values within , and further strengthen the tension with early universe measurements of the expansion rate of the Universe.

### v2

@@ 4.2 The global slope and zero-points of anchor PLRs
The universality of the slope of Mira PLR in F160W filters is the basic assumption for the distance scale application. C. D. Huang et al. (2018) used two different slopes of mag/dex based on W. Yuan et al. (2017) LMC sample, and mag/dex based on the transformed HST F160W filter. The slopes of the globular cluster and LMC PLRs in the HST F160W filter are significantly different, presumably due to different period distribution of Miras in these systems. With three independent anchors available, we can determine a global slope of Mira PLRs in calibrated F160W magnitudes. For the LMC, we adopt the precise geometric distance to the LMC (G. Pietrzyński et al., 2019). For the NGC 4258, a geometric maser distance of mag (D. W. Pesce et al., 2020) was adopted that is precise to .
@@
To account for the possible metallicity effects, we adopt a two-parameter solution simultaneously deriving a zero-point and a metallicity coefficient. For the cluster Miras, we derive a mean metallicity of with a scatter of dex after excluding the most metal-poor cluster. For the LMC, a metallicity of dex is taken based on age-metallicity relations for intermediate-age stellar populations in the field and cluster stars (W. Narloch et al., 2023). M. W. Feast (1996) also suggested [Fe/H] dex for Miras in the LMC based on a comparison with cluster variables. Given the lack of metallicity measurements for stellar populations in NGC 4258, we adopt a value of dex based on HII region emission line flux measurements from W. Yuan et al. (2022). This value is consistent with stellar metallicity of blue supergiants (R.-P. Kudritzki et al., 2024) and has been used in the Cepheid-based distance ladder (A. G. Riess et al., 2022a). While the adopted metallicity is appropriate for younger stellar populations, considering no significant metallicity gradient in NGC 4258 (W. Yuan et al., 2022; R.-P. Kudritzki et al., 2024) and the lack of stellar metallicity measurements, this approximation is adopted with a relatively larger uncertainty of 0.15 dex. For the Milky Way (MW) field stars, we assume solar metallicity for Miras, but this was not included in the quantitative analysis. Using the aforementioned fixed slope, we estimated the following global solution:
where the mean metallicity of the anchor sample is [Fe/H] dex. We find a small positive metallicity coefficient such that metal-rich Miras are fainter. However, this quantification of metallicity effect is based on our approximation of mean metallicities of Miras in the LMC, in globular clusters and NGC 4258. Fig. 6 displays this linear fit to the zero-points of three anchor galaxies as a function of their adopted metallicities. We find a small variation in the zero-point of mag with or without the metallicity term, further suggesting a small effect in the luminosity due to variations in iron abundance.
Furthermore, no metallicity measurements of Miras or intermediate-age populations are available in SNe Ia host galaxies. C. D. Huang et al. (2024) used an oxygen abundance gradient based-on HII regions to derive an approximate metallicity range for M101 Miras that falls between solar and LMC values. We assume that the SNe Ia host galaxies have metallicities similar to NGC 4258 with a possible difference of dex, which may increase up to dex with respect to LMC Miras. Given the approximations involved in determining the metallicity coefficient, we do not include this term in the subsequent analysis, but add a systematic uncertainty term accounting for the difference in mean metallicities between anchor and host galaxies.
@@ 4.3 Distances to Type Ia Supernovae host galaxies
The second step of the distance ladder involves using primary calibration of stellar standard candles to determine distances to SNe Ia host galaxies and calibrate the peak fiducial luminosity of SNe Ia. For the Mira distance scale, there are two supernovae host galaxies observed with HST — SN 2005df in NGC 1559 (C. D. Huang et al., 2020) and SN 2011fe in M101 (C. D. Huang et al., 2024). We used a sample of 110 Miras in NGC 1559 and 211 Miras in M101 from the aforementioned studies along with their periods and calibrated F160W magnitudes including corrections due to crowding effects. We applied corrections for the Milky Way foreground extinction as described in C. D. Huang et al. (2024).
@@ 4.4 Hubble flow, , and systematics
Table 6 provides a detailed breakdown of the error budget on the . The total uncertainty is dominated by the statistical uncertainty on the fiducial peak luminosity of SNe Ia in two host galaxies. With a global slope that is determined using Miras in three-anchor galaxies, we obtained an improved constraint on the anchor zero-point luminosity of a Mira with d period. The absolute value of the anchor zero-point is similar to C. D. Huang et al. (2024) value, but the uncertainty is reduced significantly considering that the metallicity systematics are also included in the anchor zero-point. The metallicity systematic is propagated through the coefficient of with a plausible difference in the mean metallicity of dex between NGC 4258/Cluster/LMC and SNe Ia host galaxies. We also include a systematic error of anchor distances and also add 0.02 mag for the to F160W transformations for the ground-based LMC and Cluster calibrations. The statistical uncertainties due to SNe Ia host galaxies and the slope of the PLR are similar to those from C. D. Huang et al. (2024). The uncertainty due to interstellar extinction in SNe Ia host galaxies is propagated assuming that the reddening is similar to the LMC, where a difference of 0.02 mag is obtained with and without interstellar reddening correction. The global slope together with improved anchor zero-points provides better constraints on the for individual SNe Ia host galaxies. The difference in of mag between two SNe host galaxies suggests excellent consistency in the fiducial peak luminosity of their SNe Ia. Finally, the inverse-weighted average value of fiducial peak luminosity of SNe Ia is better constrained with an error smaller than the estimated uncertainty in C. D. Huang et al. (2024). These small improvements resulting from an additional anchor in the solution provide a significant reduction in the error budget to in the Mira distance ladder.
@@ 5 Discussion
Another approach to quantifying metallicity effects involves comparing the calibrated zero-points of the PLRs in anchor galaxies with different metallicities. Using cluster, LMC, and NGC 4258 Miras, we find a small slope ( mag/dex) of the variation in the zero-points as a function of metallicity. This determination of metallicity effects is also very sensitive to the uncertainties on distances to these systems, as well as the adopted mean metallicities for their Mira stars. For the distant systems observed with HST, we assume the mean metallicity based on HII regions that is used in the Cepheid distance ladder (A. G. Riess et al., 2022b). This approximation may not be appropriate for intermediate-age Mira populations, and therefore, we do not include the metallicity term in the distance scale analysis. However, these mean metallicity values can be used to estimate the relative difference in abundances between anchor and host galaxies, assuming that it is similar for young and intermediate-age populations. This allowed us to provide a reasonable estimate of systematic uncertainty due to metallicity in the determinations. If we propagate a metallicity term of mag/dex in the Mira distance ladder adopting the metallicity of SNe Ia host galaxies from A. G. Riess et al. (2022a), we find a small difference of units in . Homogeneous spectroscopic metallicity determinations of intermediate-age stars or AGB stars in these galaxies will be useful to precisely quantify the metallicity effects based on the zero-points of anchor PLRs.
@@
Finally, the Mira based km s-1 Mpc-1 derived in this work provides an independent comparison to Cepheid and TRGB based determinations. As previously mentioned, the uncertainty on the value primarily comes from the peak fiducial luminosity of SNe Ia, and additional SNe Ia hosts with Mira variables are required to minimize these statistical uncertainties. Nevertheless, our accurate determination of is in excellent agreement with the three-anchor baseline solution from Cepheid-based local measurements. The value based on intermediate-age Mira stars is also in agreement with Population II TRGB-based measurements from W. L. Freedman et al. (2025). Higher-precision Mira-based values may provide insight into possible systematics due to stellar standard candles of different age and metallicity populations. Currently, our estimates are in excellent agreement with Cepheid-based values within , and further strengthen the tension with early universe measurements of the expansion rate of the Universe.